OpenStudy (reese.jordan):

PLEASE HELP!!!!!! two chords ab and cd are equidistant from the center of a circle AB=3x+11 and CD=39-x. Solve for x

OpenStudy (luckycoins888):

AB = 3x + 11 = CD = 39 - x 4x = 28 x = 7

OpenStudy (reese.jordan):

thank you!!